OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL Which are The rational roots for f(x)=5x^2-3x+3 A +/-1 B +/-3 C +/-1/3 D +/-5 E +/-1/5

OpenStudy (anonymous):

Choose all correct answers

OpenStudy (noelgreco):

Just by seeing that the discriminant is < 0 it is apparent that there are no real solutions. All of the solutions you propose are real.

OpenStudy (anonymous):

As NoelGreco said, there are no real solutions to that equation, so none of those options are correct. a = 5 b = -3 c = 3, so if you were to plug a,b,c into the quadratic formula it wouldn't work because the discriminant (the b^2 - 4ac in the quadratic formula: what is under the root sign) cannot be negative... since you can't take the square root of a negative number and get a real solution. b^2 - 4ac = = (-3)^2 - 4(5)(3) = 9 - 60 = (-) The discrimant is negative in this case so there are no real solutions.